help me solve 3 to the 12 power and 27 to the 8 power
1
Expert's answer
2012-10-04T08:35:50-0400
3^12 = 3^6 * 3^6 = 9^5 * 9^5 = ( 9^2 * 9^3 )² = ( 81 * 729 )² = 59049² = 3486784401.

27^8 = ( 27² )^4 = 729^4 = (729²)² = (531441)² = 282429536481.